### v1

@@ 4.1 Properties of contaminating unmasked sources
To estimate the flux distribution of the contaminating signal, we construct $L_{\mathrm{X}}-M_{*}$ for three types of unmasked X-ray sources. Meanwhile, we also investigate the spectral properties of these to be used in count rate conversion in spectral fitting. The details of the modeling are given in the subsections below for each component.
@@ 4.1.1 Unresolved haloes
We use the X-ray halo $L_{\mathrm{X}}-M_{*}$ relation from RASS galaxy stacking analysis (Anderson et al. 2015) to predict the X-ray luminosity of bound gas in unmasked galaxy (group and cluster) halos. Because we exclude eRASS:4 detected X-ray sources and clusters detected in the optical band with richness higher than 20 before stacking, to obtain the $L_{\mathrm{X}}-M_{*}$ relation that meets the same selection, we use our galaxy cluster mock catalog that is constructed for eRASS predictions. The catalog is based on HugeMDPL dark matter only N-body simulation (Klypin et al. 2016), and the X-ray properties of each halo are painted by a phenomenological modeling approach (Comparat et al. 2020) and then scaled to agree the observed $L_{\mathrm{X}}-M_{*}$ relation from RASS galaxy stacking results of Anderson et al. (2015). By applying the same source selection criteria, we tabulate the averaged rest-frame $L_{\mathrm{0.5-2keV}}-M_{*}$ relation for unmasked X-ray halos in different redshift bins.
@@
The galaxy sample used in Anderson et al. (2015) is selected from an SDSS-DR7 based New York University Value Added Galaxy Catalogue (NYU-VAGC) (Blanton et al. 2005). The stellar mass in NYU-VAGC is calculated based on the method of Blanton & Roweis (2007). It is known that the Blanton & Roweis (2007) stellar mass is systematically lower than many other estimates at the high stellar mass end (e.g. Li & White 2009; Bernardi et al. 2010; Moustakas et al. 2013). Because the X-ray halo $L_{\mathrm{X}}-M_{*}$ relation has a steep slope of $\gtrsim 3$, a small underestimate of the stellar mass at the high mass end could introduce a huge difference on the estimated unmasked X-ray halo emission. Therefore, we need to properly modify the current scaling relation to match the Zou et al. (2019) stellar mass scheme. We selected a $z<0.2$ subsample of Zou et al. (2019) galaxies with available spectroscopic redshift to cross-match the NYU-VAGC catalog and obtain the discrepancy between the two stellar mass schemes. We use this stellar mass discrepancy to empirically modify the tabulated $L_{\mathrm{0.5-2keV}}-M_{*}$ relation. The modified X-ray halo $L_{\mathrm{X}}-M_{*}$ relation is plotted in the top left panel of Fig. 6).
@@
To convert the rest-frame 0.5–2 keV luminosity to the 0.3–1.2 keV observed count rate, the spectral properties of the sources are necessary. For this purpose, we assume that each source can be represented by a single temperature collisional ionization equilibrium (CIE) model and estimate the temperature using the luminosity - temperature ($L_{\mathrm{X}}-kT$ ) scaling relation constructed using eFEDS galaxy cluster and groups (Bahar et al. 2022).
@@
With the estimated temperature of each source in the mock catalog by using the $L_{\mathrm{X}}-kT$ scaling relation, we also predict a synthesis spectral model of unmasked X-ray halos for spectral analysis. After applying the selection, we construct the multi-temperature differential emission measure (DEM) models for each redshift bin (left panel of Fig. 17). The spectral model on each single temperature is calculated using Astrophysical Plasma Emission Code (APEC) with $Z=0.4Z_{\sun}$. Then we use the redshift and length distributions of the selected filament sample to create the synthesis spectral model as,
@@
where $l_{i}$ is the physical length of the $i$-th filament, $z_{i}$ is the redshift of the $i$-th filament, $F(E,z_{i})$ is the multi-temperature differential emission measure spectrum at the redshift of $z_{i}$, $D_{\mathrm{L}}$ is the luminosity distance and is a function of redshift. The constructed model spectrum is shown in the right panel of Fig. 17.
@@ 4.1.2 Unresolved AGN
We calculate the galaxy averaged $L_{\mathrm{X}}-M_{*}$ relation for unresolved AGN using an AGN mock catalog based on the high-resolution Uchuu dark matter only N-body simulation (Ishiyama et al. 2021). The AGN properties are painted to dark matter halos based on an empirical method (see Comparat et al. 2019, for further details), where the AGN model is based on the observed AGN 2–10 keV X-ray luminosity function (Aird et al. 2015), duty cycle (Georgakakis et al. 2017), obscuration distribution (Ricci et al. 2017), spectral model (Liu et al. 2017), and parameters of AGN halo occupation distribution (Comparat et al. 2023).
We apply a selection criteria that removes AGN with $F_{\mathrm{0.5-2keV}}>10^{-14}$ erg s-1 cm-2, which would be detected in eRASS:4, to select sources in the mock catalog. In each redshift-stellar mass bin, we sum the rest-frame 0.5–2 keV luminosity and divide it by the total number of galaxies in the same bin to obtain the galaxy averaged unresolved AGN luminosity. In our method, a galaxy-averaged scaling relation will distribute the unmasked AGN flux to all galaxies. Therefore, we do not need to additionally predict whether a galaxy hosts an AGN. The scaling relations at different redshifts are shown in the top right panel of Fig. 6.
@@ 4.1.3 XRB in galaxies
We use a stellar mass and star forming rate (SFR) dependent model of integrated X-ray luminosity of XRB constrained by Chandra observations of local galaxies (Lehmer et al. 2019), and the UniverseMachine model of redshift dependent SMF and SFR (Behroozi et al. 2019) to construct the averaged $L_{\mathrm{X}}-M_{*}$ relations of XRB at different redshifts. The results are plotted in the bottom left panel of Fig. 6, where the relations show redshift evolution at $M_{*}<10^{11}$ $M_{\odot}$, which is the result of the redshift evolution of star-forming rate and star-forming galaxy fraction.
We adopt a power-law distribution with intrinsic absorption spectral model for flux conversion, with parameters $\Gamma=1.7$ and $n_{\mathrm{H}}=2\times 10^{21}$ cm-2 suggested by observations (Lehmer et al. 2019).
@@ 4.1.4 Comparison with observed $L_{\mathrm{X}}-M_{*}$ relations
Recently, Zhang et al. (2024) investigated the $L_{\mathrm{X}}-M_{*}$ scaling relation of galaxies using eRASS:4 data. We compare our constructed total unresolved galactic X-ray sources and their $L_{\mathrm{X}}-M_{*}$ relations with their results from eROSITA observations. We note that the stellar mass used in Zhang et al. (2024) is adopted from Tinker (2021), which was originally estimated by Chen et al. (2012). Similar to the analysis in Sect. 4.1.1, we scale the observed $L_{\mathrm{X}}-M_{*}$ scaling relation in Zhang et al. (2024) to match the stellar mass scheme of Zou et al. (2019) for comparison. In the bottom right panel of Fig. 6, we plot the total unresolved source scaling relation model together with that of a central galaxy sample from Zhang et al. (2024), whose median redshift is $\sim 0.1$. Our constructed relation at $z=0.2$ has the same slope at different stellar masses as the the observed relation, and the slight luminosity difference meets the expected redshift revolution of AGN and XRB scaling relations.
@@ 4.2 Fractions in stacked signals
Next, we calculate the integrated signal of unmasked halos, AGNs and XRBs in the 0–10 Mpc transverse distance range to account for the contamination in our imaging analysis. The total contribution is $37\%$, with each of $20\%$, $12\%$, and $5\%$ for unmasked halos, AGNs and XRBs, respectively (also see Fig. 7). We further divide the stellar mass range into two subranges of $119.5<\log M_{*}(M_{\sun})<$ 11 and $1211<\log M_{*}(M_{\sun})<$ 12, and repeat the same analysis. The results of fractions are listed in Table 2. We find that in the high stellar mass bin, where the unmasked galaxies are mostly central galaxies of group-size halos, the unmasked signal is dominated by halos gas. Whereas in the low stellar mass bin, unmasked AGN and XRB have more contribution. We use a bootstrapping method to estimate the statistical uncertainties of the fraction. In short, we randomly select galaxies with replacement to construct bootstrapping samples, and calculate the standard deviations of the fractions estimated using the bootstrapping samples as the statistical uncertainties. The relative uncertainties are $6.9\%$, $1.8\%$ and $1.4\%$ for unmasked halos, AGNs and XRB, respectively. Given that the unmasked halos contribute about half of the unmasked signal, the averaged statistical uncertainty is below $5\%$.
Because the uncertainty on photometric redshifts that we used in this analysis is not negligible and is a function of magnitude (it can reach a few tens percent for faint galaxies Zou et al. 2019), it has a potential impact on the recovered unresolved contamination signal. The redshift range of the filament sample we select is $0.60.2<z<$ 0.6, and that of the galaxies we used to model unmasked source contribution is $0.70.15<z<$ 0.7. It means that for a galaxy whose true redshift $z_{\mathrm{true}}$ is within the filament redshift range, but the measured redshift $z_{\mathrm{phot}}$ is outside the redshift range of the selected galaxies, it may not be taken into account in the unmasked source profile. On the contrary, for a galaxy with $z_{\mathrm{true}}$ outside of the filament redshift range but $z_{\mathrm{phot}}$ within that range, because it is not spatially correlated with the filament we selected, it may not contribute to the unmasked source profile either. Therefore, we note that the fraction of unmasked source signals could be underestimated in our analysis due to the galaxy photo-z uncertainties.
We use a Monte-Carlo method to calculate the recovered flux fraction of galaxies at a given redshift with given redshift uncertainties. For creating the randomized redshifts, we follow the definition of normalized redshift error used by Zou et al. (2019), $\Delta z_{\mathrm{norm}}\equiv(z_{\mathrm{phot}}-z_{\mathrm{spec}})/(1+z_{\mathrm{spec}})$, and assume that $z_{\mathrm{spec}}=z_{\mathrm{true}}$. The ratio between the recovered flux and the true flux of a galaxy is
@@
where $n_{\mathrm{MC}}$ is the number of realizations, $z_{i}$ is the $i$ th randomized redshift value, the value of $w_{i}$ is one if $0.70.15<z_{i}<$ 0.7 and is zero elsewhere. The results are shown as dashed lines in Fig. 8. Based on the $90\%$ percentile of the $r$-band magnitude distribution of the galaxies we selected, and the $r-\Delta z$ relation shown in Fig. 5 of Zou et al. (2019), we predict the recovered flux fractions for unmasked galaxies in the $119.5<\log M_{*}(M_{\sun})<$ 11 and $1211<\log M_{*}(M_{\sun})<$ 12 stellar mass ranges, which are shown in the crossed and triangle curve in Fig. 8, respectively. The recovered fraction of galaxies in the high stellar bin is around one, which is due to the high quality of photometric redshift of bright central galaxies. This fraction is $\sim 80\%$ in the low stellar mass bin when $z>0.3$. It means that even if we take the impact of photo-z uncertainty into account, the total contribution from unresolved sources can be as high as $\sim 40\%$.
The $\sim 40\%$ fraction of the unmasked source contribution suggests that the remaining $\sim 60\%$ of the stacked signal is from the WHIM in the filaments, which reach a $5.4\sigma$ detection significance. We note that this fraction relies on the galaxy X-ray contamination modeling, which is based on our current knowledge of different X-ray emitters in galaxy systems at different scales. Future better understanding of the all different types of X-ray sources in the Universe will bring us more accurate results of the contamination fraction.
Figure 7: The contribution of unmasked X-ray halos (green), AGNs (light blue), and XRBs (grey) to the total stacked signal (orange). To find the net WHIM signal, the contribution from these contaminating signals is modeled out in the analysis. The signal from WHIM is $\sim 60\%$ based on the unmasked signal estimation.
@@
Figure 8: Fraction of recovered flux of unresolved sources as a function of true redshift (x-axis) and normalized redshift uncertainty (dashed lines in color). The triangle and crossed lines are the fractions for our selected galaxies in the $1211<\log M_{*}(M_{\sun})<$ 12 and $119.5<\log M_{*}(M_{\sun})<$ 11 stellar mass bins, respectively. For high stellar mass galaxies that contribute most of the unmasked signals, the impact of photometric redshift uncertainty is only a few percent.
@@ 5 Spectral stacking
In this section, we describe the X-ray spectral analysis of the filaments. Due to the faint nature of the cosmic filaments, to increase the signal-to-noise in the data, we stack the spectra extracted from individual filaments by following the recipe provided in Bulbul et al. (2014). The method successfully produces the stacked rest-frame spectrum of filaments at different redshifts. We study the spectral properties and constrain the temperature of the detected WHIM using the stacked filament spectra through this analysis.

### v2

@@ 4.1 Properties of contaminating unmasked sources
To estimate the flux distribution of the contaminating signal, we constructed $L_{\mathrm{X}}-M_{*}$ for three types of unmasked X-ray sources. Meanwhile, we also investigated the spectral properties to be used in count rate conversion in spectral fitting. The details of the modeling for each component are given in the subsections below.
@@ 4.1.1 Unresolved haloes
We used the X-ray halo $L_{\mathrm{X}}-M_{*}$ relation from RASS galaxy stacking analysis (Anderson et al. 2015) to predict the X-ray luminosity of bound gas in unmasked galaxy (group and cluster) halos. Because we excluded eRASS:4-detected X-ray sources and clusters detected in the optical band with a richness higher than 20 before stacking, to obtain the $L_{\mathrm{X}}-M_{*}$ relation that meets the same selection – we use our galaxy cluster mock catalog that is constructed for eRASS predictions. The catalog is based on HugeMDPL dark matter-only N-body simulation (Klypin et al. 2016), while the X-ray properties of each halo are illustrated by a phenomenological modeling approach (Comparat et al. 2020) and then scaled to agree with the observed $L_{\mathrm{X}}-M_{*}$ relation from RASS galaxy stacking results of Anderson et al. (2015). We tabulated the averaged rest-frame $L_{\mathrm{0.5-2keV}}-M_{*}$ relation for unmasked X-ray halos in different redshift bins by applying the same source selection criteria.
@@
The galaxy sample used in Anderson et al. (2015) was selected from an SDSS-DR7 based New York University Value Added Galaxy Catalogue (NYU-VAGC) (Blanton et al. 2005). The stellar mass in NYU-VAGC is calculated based on the method of Blanton & Roweis (2007). It is known that the Blanton & Roweis (2007) stellar mass is systematically lower than many other estimates at the high stellar mass end (e.g., Li & White 2009; Bernardi et al. 2010; Moustakas et al. 2013). Because the X-ray halo $L_{\mathrm{X}}-M_{*}$ relation has a steep slope of $\gtrsim 3$, a small underestimate of the stellar mass at the high-mass end could introduce a huge difference on the estimated unmasked X-ray halo emission. Therefore, we need to properly modify the current scaling relation to match the Zou et al. (2019) stellar mass scheme. We selected a $z<0.2$ subsample of Zou et al. (2019) galaxies with available spectroscopic redshift to cross-match the NYU-VAGC catalog and obtained the discrepancy between the two stellar mass schemes. We used this stellar mass discrepancy to empirically modify the tabulated $L_{\mathrm{0.5-2keV}}-M_{*}$ relation. The modified X-ray halo $L_{\mathrm{X}}-M_{*}$ relation is plotted in the top left panel of Fig. 6).
@@
To convert the rest-frame 0.5–2 keV luminosity to the 0.3–1.2 keV observed count rate, the spectral properties of the sources are needed. For this purpose, we assumed that each source can be represented by a single temperature collisional ionization equilibrium (CIE) model and we estimated the temperature using the luminosity - temperature ($L_{\mathrm{X}}-kT$ ) scaling relation constructed using eFEDS galaxy cluster and groups (Bahar et al. 2022).
@@
With the estimated temperature of each source in the mock catalog by using the $L_{\mathrm{X}}-kT$ scaling relation, we were also able to predict a synthesis spectral model of unmasked X-ray halos for spectral analysis. After applying the selection, we construct the multi-temperature differential emission measure (DEM) models for each redshift bin (left panel of Fig. 17). The spectral model on each single temperature is calculated using Astrophysical Plasma Emission Code (APEC) with $Z=0.4Z_{\sun}$. Then, we used the redshift and length distributions of the selected filament sample to create the synthesis spectral model as:
@@
where $l_{i}$ is the physical length of the $i$-th filament, $z_{i}$ is the redshift of the $i$-th filament, $F(E,z_{i})$ is the multi-temperature differential emission measure spectrum at the redshift of $z_{i}$, and $D_{\mathrm{L}}$ is the luminosity distance as a function of redshift. The constructed model spectrum is shown in the right panel of Fig. 17.
@@ 4.1.2 Unresolved AGN
We calculate the galaxy averaged $L_{\mathrm{X}}-M_{*}$ relation for unresolved AGNs using an AGN mock catalog based on the high-resolution Uchuu dark matter only N-body simulation (Ishiyama et al. 2021). The AGN properties are painted to dark matter halos based on an empirical method (see Comparat et al. 2019, for further details), where the AGN model is based on the observed AGN 2–10 keV X-ray luminosity function (Aird et al. 2015), duty cycle (Georgakakis et al. 2017), obscuration distribution (Ricci et al. 2017), spectral model (Liu et al. 2017), and parameters of AGN halo occupation distribution (Comparat et al. 2023).
We applied a selection criteria that removes AGNs with $F_{\mathrm{0.5-2keV}}>10^{-14}$ erg s-1 cm-2, which would be detected in eRASS:4, to select sources in the mock catalog. In each redshift-stellar mass bin, we summed the rest-frame 0.5–2 keV luminosity and divide it by the total number of galaxies in the same bin to obtain the galaxy averaged unresolved AGN luminosity. In our method, a galaxy-averaged scaling relation would distribute the unmasked AGN flux to all galaxies. Therefore, we do not need to additionally predict whether a galaxy hosts an AGN. The scaling relations at different redshifts are shown in the top right panel of Fig. 6.
@@ 4.1.3 XRB in galaxies
We used an integrated X-ray luminosity model that is a function of stellar mass and star forming rate (SFR) to construct the averaged $L_{\mathrm{X}}-M_{*}$ relations of XRB at different redshifts. The model is constrained by Chandra observations of local galaxies (Lehmer et al. 2019), and the UniverseMachine model of redshift dependent SMF and SFR (Behroozi et al. 2019). The results are plotted in the bottom left panel of Fig. 6, where the relations show redshift evolution at $M_{*}<10^{11}$ $M_{\odot}$ (the result of the redshift evolution of star-forming rate and star-forming galaxy fraction).
We adopted a power-law distribution with intrinsic absorption spectral model for flux conversion, with parameters of $\Gamma=1.7$ and $n_{\mathrm{H}}=2\times 10^{21}$ cm-2, as suggested by observations (Lehmer et al. 2019).
@@ 4.1.4 Comparison with observed $L_{\mathrm{X}}-M_{*}$ relations
Zhang et al. (2024) investigated the $L_{\mathrm{X}}-M_{*}$ scaling relation of galaxies using eRASS:4 data. We compared our constructed total unresolved galactic X-ray sources and their $L_{\mathrm{X}}-M_{*}$ relations with their results from eROSITA observations. We note that the stellar mass used in Zhang et al. (2024) was adopted from Tinker (2021), which was originally estimated by Chen et al. (2012). Similarly to the analysis in Sect. 4.1.1, we scaled the observed $L_{\mathrm{X}}-M_{*}$ scaling relation in Zhang et al. (2024) to match the stellar mass scheme of Zou et al. (2019) for comparison. In the bottom right panel of Fig. 6, we plot the total unresolved source scaling relation model together with that of a central galaxy sample from Zhang et al. (2024), whose median redshift is $\sim 0.1$. Our constructed relation at $z=0.2$ has the same slope at different stellar masses as the observed relation and the slight luminosity difference meets the expected redshift revolution of AGN and XRB scaling relations.
@@ 4.2 Fractions in stacked signals
Next, we calculated the integrated signal of unmasked halos, AGNs, and XRBs in the 0–10 Mpc transverse distance range to account for the contamination in our imaging analysis. The total contribution is $37\%$, with each of $20\%$, $12\%$, and $5\%$ for unmasked halos, AGNs, and XRBs, respectively (also see Fig. 7). We further divided the stellar mass range into two subranges ($119.5<\log M_{*}(M_{\sun})<$ 11 and $1211<\log M_{*}(M_{\sun})<$ 12) and repeated the same analysis. The results of the fractions are listed in Table 2. In the high stellar mass bin, where the unmasked galaxies are mostly central galaxies of group-size halos, the unmasked signal is dominated by halos gas; whereas in the low stellar mass bin, unmasked AGNs and XRBs demonstrate more of a contribution. We used a bootstrapping method to estimate the statistical uncertainties of the fraction. In short, we randomly select galaxies with replacements to construct bootstrapping samples and calculated the standard deviations of the fractions estimated using the bootstrapping samples as the statistical uncertainties. The relative uncertainties are $6.9\%$, $1.8\%$, and $1.4\%$ for unmasked halos, AGNs, and XRB, respectively. Given that the unmasked halos contribute about half of the unmasked signal, the averaged statistical uncertainty is below $5\%$.
Because the uncertainty on photometric redshifts that we used in this analysis is not negligible and is a function of magnitude (it can reach up to a few tens of percent for faint galaxies Zou et al. 2019), it has a potential impact on the recovered unresolved contamination signal. The redshift range of the filament sample we select is $0.60.2<z<$ 0.6, while the range for the galaxies we used to model unmasked source contribution is $0.70.15<z<$ 0.7. This means that for a galaxy whose true redshift $z_{\mathrm{true}}$ is within the filament redshift range, but the measured redshift $z_{\mathrm{phot}}$ is outside the redshift range of the selected galaxies, it may not be taken into account in the unmasked source profile. On the contrary, for a galaxy with $z_{\mathrm{true}}$ outside of the filament redshift range but $z_{\mathrm{phot}}$ within that range, because it is not spatially correlated with the filament we selected, it may not contribute to the unmasked source profile either. Therefore, we note that the fraction of unmasked source signals could be underestimated in our analysis due to the galaxy photo-z uncertainties.
We use a Monte Carlo method to calculate the recovered flux fraction of galaxies at a given redshift with given redshift uncertainties. For creating the randomized redshifts, we follow the definition of normalized redshift error used by Zou et al. (2019), $\Delta z_{\mathrm{norm}}\equiv(z_{\mathrm{phot}}-z_{\mathrm{spec}})/(1+z_{\mathrm{spec}})$ and assume that $z_{\mathrm{spec}}=z_{\mathrm{true}}$. The ratio between the recovered flux and the true flux of a galaxy is
@@
where $n_{\mathrm{MC}}$ is the number of realizations, and $z_{i}$ is the $i$ th randomized redshift value, while the value of $w_{i}$ is 1 if $0.70.15<z_{i}<$ 0.7 and it is 0 otherwise. The results are shown as dashed lines in Fig. 8. Based on the $90\%$ percentile of the $r$-band magnitude distribution of the galaxies we selected and the $r-\Delta z$ relation shown in Fig. 5 of Zou et al. (2019), we were able to predict the recovered flux fractions for unmasked galaxies in the $119.5<\log M_{*}(M_{\sun})<$ 11 and $1211<\log M_{*}(M_{\sun})<$ 12 stellar mass ranges, which are shown in the crossed and triangle curve in Fig. 8, respectively. The recovered fraction of galaxies in the high stellar bin is around one due to the high quality of photometric redshift of bright central galaxies. This fraction is $\sim 80\%$ in the low stellar mass bin when $z>0.3$. It means that even if we take the impact of photo-z uncertainty into account, the total contribution from unresolved sources can be as high as $\sim 40\%$.
The $\sim 40\%$ fraction of the unmasked source contribution suggests that the remaining $\sim 60\%$ of the stacked signal is from the WHIM in the filaments, reaching a $5.4\sigma$ detection significance. This fraction relies on the galaxy X-ray contamination modeling, which is based on our current knowledge of different X-ray emitters in galaxy systems at different scales. A better understanding of all the different types of X-ray sources in the Universe will offer more accurate results of the contamination fraction in the future.
Figure 7: Contribution of unmasked X-ray halos (green), AGNs (light blue), and XRBs (grey) to the total stacked signal (orange). To find the net WHIM signal, the contribution from these contaminating signals is modeled out in the analysis. The signal from WHIM is $\sim 60\%$ based on the unmasked signal estimation.
@@
Figure 8: Fraction of recovered flux of unresolved sources as a function of true redshift (x-axis) and normalized redshift uncertainty (dashed lines in color). The triangle and crossed lines are the fractions for our selected galaxies in the $1211<\log M_{*}(M_{\sun})<$ 12 and $119.5<\log M_{*}(M_{\sun})<$ 11 stellar mass bins, respectively. The impact of photometric redshift uncertainty is only a few percent for high stellar mass galaxies that contribute most of the unmasked signals.
@@ 5 Spectral stacking
In this section, we describe the X-ray spectral analysis of the filaments. Due to the faint nature of the cosmic filaments, to increase the signal-to-noise ratio (S/N) in the data, we stacked the spectra extracted from individual filaments by following the recipe provided in Bulbul et al. (2014). The method successfully produces the stacked rest-frame spectrum of filaments at different redshifts. Through this analysis, we study the spectral properties and constrain the temperature of the detected WHIM using the stacked filament spectra.